# Subspace State Transfer in Coined Quantum Walks

### v1

@@ Abstract
We study a transport phenomenon in certain coined quantum walks where a subspace of states localized at a vertex gets transferred to another vertex. We first develop characterizations for perfect and pretty good subspace state transfer using the spectral properties of a Hermitian weighted digraph obtained from the underlying graph. We then provide a polynomial-time algorithm that tests whether perfect pointwise subspace state transfer occurs at an integer step, given that the subspace and coins are rational. Finally, we construct several infinite families of examples that admit perfect pointwise $d$-dimensional subspace state transfer where $d\geq 2$.
@@ 1 Introduction
Switching to coins that reflect about higher-dimensional subspaces brings several advantages. For example, as shown by Kubota and Segawa [18], the only $4$-regular circulant graphs that admit perfect state transfer in Grover coined walks are $X({\mathbb{Z}}_{2m},\pm\{a,m-a\})$ where $m$ is odd or oddly even, and the minimum transfer time is at least $m$. On the other hand, by Theorem 9.3 of this paper, assigning negative Grover coins to the sender and receiver vertices enables perfect state transfer on more circulant graphs ($X({\mathbb{Z}}_{2m},\pm\{a,m-a\})$ for any $m$), increases the dimension of subspaces of states transferred, and reduces the transfer time to a constant.
@@
Our characterization connects subspace state transfer to properties of certain Hermitian weighted digraph. This leds to a polynomial-time algorithm in that tests whether perfect pointwise subspace state transfer occurs at an integer step, given that the subspace and coins are rational. By marking vertices in Grover coined walks, we construct several infinite families of examples that admit perfect pointwise subspace state transfer.
@@ 2 Reflection coins and weighted digraphs
for some matrix $N_{u}$ with orthonormal columns. The rows of $N_{u}$ are indexed by the outgoing arcs of $u$, and the columns of $N_{u}$, if any, can be seens as indexed by some clones of $u$. We can then write
@@
is some weighted incidence matrix of the arcs of $X$ and the clones of vertices in $X$. The spectrum of $U=RC$ is closely related to the matrix $H=N^{*}RN$. Here $H$ also carries some combinatorial meaning: we can think of it as the Hermitian adjacency matrix of some digraph $Y$ constructed from $X$, where the vertices of $Y$ are the clones of $X$, and the weights on the arcs between these clones are determined by the entries of $N$. Figure 5 illustrates some reflection coin operators on the Peterson graph and their associated Hermitian weighted digraphs. We will describe these digraphs in more detail in Section 3.
@@
The following spectral relation between $U$ and $H$ is a consequence of [11, Chapter 2]. We omit the proof as it is analogous to that of [7, Theorem 3.3].
@@
[11, Chapter 2] Let $U=R(2NN^{*}-I)$ be the transition matrix where $N$ has orthonormal columns. Let $H=N^{*}RN$. The eigenvalues of $U$ are $1$, 1-1 and $e^{\pm i\theta}$, where $\theta=\arccos\lambda$ for some eigenvalue $\lambda\in(-1,1)$ of $H$. Moreover, if $F_{\theta}$ denotes the $e^{i\theta}$-eigenprojection of $U$, and $E_{\lambda}$ denotes the $\lambda$-eigenprojection of $H$, then $F_{\theta}$ and $E_{\lambda}$ are related in the following way.
@@ 3 Subspace state transfer
We will also consider a relaxation of this phenomenon called pretty good subspace state transfer. That is, there exists a subspace $W$ of $\operatorname{col}(C_{a}+I)$ such that for any $\epsilon>0$ and any vector $w\in W$, there is a time $t$ such that $U^{t}x_{a}(w)$ is $\epsilon$-close to some coin state at $b$:
$\left\lVert U^{t}x_{a}(w)-x_{b}(v)\right\rVert<\epsilon$
where $v\in\operatorname{col}(C_{b}+I)$.
Any basis for $x_{a}(W)\cup x_{b}(V)$ extends to a coin basis, and relative to this basis, $\dim(W)$ clones of $a$ arise from $W$, and $\dim(V)$ clones of $b$ arise from $V$. For convenience, we will refer to such basis as a $(W,V)$-induced coin basis, and those special clones as $W$-clones of $a$ and $V$-clones of $b$, respectively.
We develop a characterization for perfect subspace state transfer using the aforementioned coin basis. For any subset $S$ of vertices in $Y$, the characteristic matrix of $S$, denoted $D_{S}$, is the diagonal matrix whose $uu$-entry is $1$ if $u\in S$, and 0 if $u\notin S$. For any two matrices $M$ and $N$ with the same size, let $\langle M,N\rangle$ denote their matrix inner product, that is,
@@
$\langle U^{t}ND_{S},ND_{T}\rangle=\langle D_{T},D_{T}\rangle.$
@@
$\left\lvert\langle U^{t}ND_{S},ND_{T}\rangle\right\rvert^{2}\leq\langle D_{S},D_{S}\rangle\langle D_{T},D_{T}\rangle=\langle D_{T},D_{T}\rangle^{2}$
and $U^{t}ND_{S}$ must be a scalar multiple of $ND_{T}$. It follows that for any $w\in W$,
$U^{t}x_{a}(w)\in\operatorname{col}(U^{t}ND_{S})\subseteq\operatorname{col}(ND_{T})=x_{a}(V).$
@@
$\left\lVert f_{t}(H)D_{S}-D_{T}\right\rVert<\epsilon,$
@@
$\left\lvert\langle f_{t}(H)D_{S},D_{T}\rangle-d\right\rvert<\epsilon,$
@@ 4 A characterization
To find necessary and sufficient conditions for subspace state transfer, we consider the inner product $\langle f_{t}(H)D_{S},D_{T}\rangle$ with $\left\lvert S\right\rvert=\left\lvert T\right\rvert$. We have
$\left\lvert\langle f_{t}(H)D_{S},D_{T}\rangle\right\rvert$ $=\left\lvert\sum_{\lambda}f_{t}(\lambda)\langle E_{\lambda}D_{S},D_{T}\rangle\right\rvert\leq\sum_{\lambda}\left\lvert f_{r}(\lambda)\right\rvert\left\lvert\langle E_{\lambda}D_{S},D_{T}\rangle\right\rvert$ (3)
$\leq\sum_{\lambda}\left\lvert\langle E_{\lambda}D_{S},D_{T}\rangle\right\rvert=\sum_{\lambda}\left\lvert\langle E_{\lambda}D_{S},E_{\lambda}D_{T}\rangle\right\rvert$ (4)
$\leq\sum_{\lambda}\sqrt{\langle E_{\lambda}D_{S},E_{\lambda}D_{S}\rangle\langle E_{\lambda}D_{T},E_{\lambda}D_{T}\rangle}$ (5)
$\leq\sqrt{\sum_{\lambda}\langle E_{\lambda}D_{S},E_{\lambda}D_{S}\rangle}\sqrt{\sum_{\lambda}\langle E_{\lambda}D_{T},E_{\lambda}D_{T}\rangle}$ (6)
$=\sqrt{\sum_{\lambda}\langle D_{S},D_{S}\rangle}\sqrt{\sum_{\lambda}\langle D_{T},D_{T}\rangle}=\sqrt{\left\lvert S\right\rvert\left\lvert T\right\rvert}=\langle D_{T},D_{T}\rangle.$
Note that the inequality in (3) is tight if and only if whenever $\langle E_{\lambda}D_{S},D_{T}\rangle$ is nonzero, it must be real, and the sign of $f_{t}(\lambda)\langle E_{\lambda}D_{S},D_{T}\rangle$ is the same for all such $\lambda$. The inequality in (4) is tight if and only if $f_{t}(\lambda)=\pm 1$. The inequality in (5) is tight if and only if $E_{\lambda}D_{S}$ is a scalar multiple of $E_{\lambda}D_{T}$. The inequality in (6) is tight if and only if $\langle E_{\lambda}D_{S},D_{S}\rangle=\langle E_{\lambda}D_{T},D_{T}\rangle$. These motivate the following definitions.
@@
$\Lambda_{S}=\{\lambda:E_{\lambda}D_{S}\neq 0\}.$
We say two subsets $S$ and $T$ of vertices in $Y$ are $\gamma$-strongly-cospectral if there is a unimodular complex number $\gamma$ such that for each eigenvalue $\lambda$ of $H$,
$E_{\lambda}D_{S}=\pm\gamma E_{\lambda}D_{T}.$
If $S$ and $T$ are $\gamma$-strongly-cospectral, then $\Lambda_{S}=\Lambda_{T}$, which can be partitioned into two disjoint sets
$\Lambda_{S,T}^{\gamma}=\{\lambda:E_{\lambda}D_{S}=\gamma E_{\lambda}D_{T}\neq 0\}$
@@
$\Lambda_{S,T}^{-\gamma}=\{\lambda:E_{\lambda}D_{S}=-\gamma E_{\lambda}D_{T}\neq 0\}.$
Sometimes $\gamma$-strong-cospectrality holds for $\gamma=\pm 1$; in this case, we simply say the two subsets are strongly-cospectral, and abbreviate $\Lambda^{\pm 1}_{S,T}$ as $\Lambda^{\pm}_{S,T}$.
@@
Let $W\subseteq\operatorname{col}(C_{a}+I)$ and $V\subseteq\operatorname{col}(C_{b}+I)$ be two subspaces with equal dimension. There is perfect $(W,V)$-transfer from $a$ to $b$ if and only if relative to some orthonormal coin basis induced by $(W,V)$, the Hermitian adjacency matrix $H$ satisfies the following.
@@
    For any eigenvalue $\lambda\in\Lambda_{ST}^{+}$, we have
    $\lambda=\cos(j\pi/t)$
    for some even integer $j$.
@@
    For any eigenvalue $\lambda\in\Lambda_{ST}^{-}$, we have
    $\lambda=\cos(j\pi/t)$
    for some odd integer $j$.
Proof. Suppose first that $f_{t}(H)D_{S}=D_{T}$. Then $\langle f_{t}(H)D_{S},D_{T}\rangle=\langle D_{T},D_{T}\rangle$. Thus inequalities (1) – (4) are all tight and $\langle f_{t}(H)D_{S},D_{T}\rangle>0$. Let $\lambda\in\Lambda_{S}$. From (3) we see that there is a scalar $\alpha_{\lambda}$ such that
$E_{\lambda}D_{S}=\alpha_{\lambda}E_{\lambda}D_{T}.$
By (4), we must have $\left\lvert\alpha_{\lambda}\right\rvert=1$. By (1), the inner product $\langle E_{\lambda}D_{S},D_{S}\rangle$ must be real with the same sign as $f_{t}(H)$. Since
$\langle E_{\lambda}D_{S},D_{S}\rangle=\alpha_{\lambda}\langle E_{\lambda}D_{T},D_{T}\rangle=\alpha_{\lambda}\langle E_{\lambda}D_{T},E_{\lambda}D_{T}\rangle,$
the scalar $\alpha_{\lambda}=1$ if $f_{t}(\lambda)>0$, and $\alpha_{\lambda}=-1$ if $f_{t}(\lambda)<0$. Finally, by (2), if $\alpha_{\lambda}=1$, then $f_{t}(\lambda)=1$, so $t\arccos\lambda$ is an even multiple of $\pi$, and if $\alpha_{\lambda}=-1$, then $f_{t}(\lambda)=-1$, so $t\arccos\lambda$ is an odd multiple $\pi$. This proves conditions (i) – (iii).
@@
$f_{t}(H)D_{S}$ $=\sum_{\lambda}f_{t}(\lambda)E_{\lambda}D_{S}$
$=\sum_{\lambda\in\Lambda_{ST}^{+}}E_{\lambda}D_{S}-\sum_{\lambda\in\Lambda_{ST}^{-}}E_{\lambda}D_{T}$
$=\sum_{\lambda\in\Lambda_{ST}^{+}}E_{\lambda}D_{T}+\sum_{\lambda\in\Lambda_{ST}^{-}}E_{\lambda}D_{T}$
$=\sum_{\lambda}E_{\lambda}D_{T}$
$=D_{T}.$
@@
We say $U$ is subspace periodic at $a$ if there is a time $t\in{\mathbb{R}}$ and a subspace $W$ of $\operatorname{col}(C_{a}+I)$ such that $x_{a}(W)$ is $U^{t}$-invariant; when $W$ is known, $U$ is said to be $W$-periodic at $a$. If, in addition, there is a unimodular complex number $\gamma$ such that
$U^{t}x_{a}(w)=\gamma x_{a}(w)$
@@
Let $W$ be a subspace of $\operatorname{col}(C_{a}+I)$. Then $U$ is pointiwse $W$-periodic at $a$ with period $t$ if and only if relative to some orthonormal coin basis induced by $(W,V)$, each eigenvalue $\lambda\in\Lambda_{S}$ of the Hermitian adjacency matrix $H$ can be written as
$\lambda=\cos(j\pi/t)$
for some even integer $j$.
@@
[20, Ch 3] Given $\alpha_{1},\alpha_{2},\ldots,\alpha_{n}\in{\mathbb{R}}$ and $\beta_{1},\beta_{2},\ldots,\beta_{n}\in{\mathbb{R}}$, the following are equivalent.
@@
$\{\langle f_{t}(H)D_{S},D_{T}\rangle:t\in{\mathbb{R}}\}$
has $d$ as a limit point. It follows from the inequalities in (5) and (6) that strong cospectrality between $S$ and $T$ is a necessary condition for pretty good $(W,V)$-transfer. Given (i), set $\sigma_{\lambda}=0$ if $\lambda\in\Lambda_{S,T}^{+}$, and $\sigma_{\lambda}=1$ if $\lambda\in\Lambda_{S,T}^{-}$. Pretty good $(W,V)$-transfer occurs from $a$ to $b$ if and only if for any $\epsilon>0$, there is a $t\in{\mathbb{R}}$ such that

### v2

@@ Abstract
We study a transport phenomenon in certain coined quantum walks where a subspace of states localized at a vertex gets transferred to another vertex. We first develop characterizations for perfect and pretty good subspace state transfer using the spectral properties of a Hermitian weighted digraph obtained from the underlying graph. We then provide a polynomial-time algorithm that tests whether pointwise perfect subspace state transfer occurs at an integer step, given that the subspace and coins are rational. Finally, we construct several infinite families of examples that admit pointwise perfect $d$-dimensional subspace state transfer where $d\geq 2$.
@@ 1 Introduction
Switching to coins that reflect about higher-dimensional subspaces brings several advantages. For example, as shown by Kubota and Segawa [21], the only $4$-regular circulant graphs that admit perfect state transfer in Grover coined walks are $X({\mathbb{Z}}_{2m},\pm\{a,m-a\})$ where $m$ is odd or oddly even, and the minimum transfer time is at least $m$. On the other hand, by Theorem 9.4 of this paper, assigning certain higher-dimensional reflection coins to the sender and receiver vertices enables perfect state transfer on more circulant graphs ($X({\mathbb{Z}}_{2m},\pm\{a,m-a\})$ for any $m$), increases the dimension of subspaces of states transferred, and reduces the transfer time to a constant.
@@
Our characterization connects subspace state transfer to properties of certain Hermitian weighted digraph. This leads to a polynomial-time algorithm that tests whether pointwise perfect subspace state transfer occurs at an integer step, given that the subspace and coins are rational. By marking vertices in Grover coined walks, we construct several infinite families of examples that admit pointwise perfect subspace state transfer.
@@ 2 Reflection coins and weighted digraphs
for some matrix $N_{u}$ with orthonormal columns. The rows of $N_{u}$ are indexed by the outgoing arcs of $u$, and the columns of $N_{u}$, if any, can be seen as indexed by some clones of $u$. We can then write
@@
is some weighted incidence matrix of the arcs of $X$ and the clones of vertices in $X$. The spectrum of $U=RC$ is closely related to that of the matrix $H=N^{*}RN$. Here $H$ also carries some combinatorial meaning: we can think of it as the Hermitian adjacency matrix of some digraph $Y$ constructed from $X$, where the vertices of $Y$ are the clones of $X$, and the weights on the arcs between these clones are determined by the entries of $N$. Figure 5 illustrates some reflection coin operators on the Peterson graph and their associated Hermitian weighted digraphs, with all weights scaled by $3$. We will describe these digraphs in more detail in Section 3.
@@
The following spectral relation between $U$ and $H$ is a consequence of [14, Chapter 2]. We omit the proof as it is analogous to that of [8, Theorem 3.3].
@@
[14, Chapter 2] Let $U=R(2NN^{*}-I)$ be the transition matrix where $N$ has orthonormal columns. Let $H=N^{*}RN$. The eigenvalues of $U$ are $1$, 1-1 and $e^{\pm i\theta}$, where $\theta=\arccos\lambda$ for some eigenvalue $\lambda\in(-1,1)$ of $H$. Moreover, if $F_{\theta}$ denotes the $e^{i\theta}$-eigenprojection of $U$, and $E_{\lambda}$ denotes the $\lambda$-eigenprojection of $H$, then $F_{\theta}$ and $E_{\lambda}$ are related in the following way.
@@ 3 Subspace state transfer
We will also consider a relaxation of this phenomenon called pretty good subspace state transfer, that is, there exists a subspace $W$ of $\operatorname{col}(C_{a}+I)$ such that for any $\epsilon>0$ and any vector $w\in W$, there is a time $t$ such that $U^{t}x_{a}(w)$ is $\epsilon$-close to some coin state at $b$:
$\left\lVert U^{t}x_{a}(w)-x_{b}(v)\right\rVert<\epsilon,\quad v\in\operatorname{col}(C_{b}+I).$
Any basis for $x_{a}(W)\cup x_{b}(V)$ extends to a coin basis, and relative to this basis, $\dim(W)$ clones of $a$ arise from $W$, and $\dim(V)$ clones of $b$ arise from $V$. For convenience, we will refer to such basis as a coin basis induced by $(W,V)$, and those special clones as $W$-clones of $a$ and $V$-clones of $b$, respectively. We develop a characterization for perfect subspace state transfer in terms of these clones. For any subset $S$ of vertices in $Y$, let $B_{S}$ be the matrix whose columns are the standard basis vectors corresponding to the vertices in $S$. For any two matrices $M$ and $N$ with the same size, let $\langle M,N\rangle$ denote their matrix inner product, that is,
@@
$\langle U^{t}NB_{S},NB_{T}\rangle=\langle B_{T},B_{T}\rangle.$
@@
$\left\lvert\langle U^{t}NB_{S},NB_{T}\rangle\right\rvert^{2}\leq\langle B_{S},B_{S}\rangle\langle B_{T},B_{T}\rangle=\langle B_{T},B_{T}\rangle^{2}$
and $U^{t}NB_{S}$ must be a scalar multiple of $NB_{T}$. It follows that for any $w\in W$,
$U^{t}x_{a}(w)\in\operatorname{col}(U^{t}NB_{S})\subseteq\operatorname{col}(NB_{T})=x_{a}(V).$
@@
$\left\lVert f_{t}(H)B_{S}-B_{T}\right\rVert<\epsilon,$
@@
$\left\lvert\langle f_{t}(H)B_{S},B_{T}\rangle-d\right\rvert<\epsilon,$
@@ 4 A characterization
To find necessary and sufficient conditions for subspace state transfer, we consider the inner product $\langle f_{t}(H)B_{S},B_{T}\rangle$ with $\left\lvert S\right\rvert=\left\lvert T\right\rvert$. We have
$\left\lvert\langle f_{t}(H)B_{S},B_{T}\rangle\right\rvert$ $=\left\lvert\sum_{\lambda}f_{t}(\lambda)\langle E_{\lambda}B_{S},B_{T}\rangle\right\rvert$
$\leq\sum_{\lambda}\left\lvert f_{t}(\lambda)\right\rvert\left\lvert\langle E_{\lambda}B_{S},B_{T}\rangle\right\rvert$ (3)
$\leq\sum_{\lambda}\left\lvert\langle E_{\lambda}B_{S},B_{T}\rangle\right\rvert$ (4)
$=\sum_{\lambda}\left\lvert\langle E_{\lambda}B_{S},E_{\lambda}B_{T}\rangle\right\rvert$
$\leq\sum_{\lambda}\sqrt{\langle E_{\lambda}B_{S},E_{\lambda}B_{S}\rangle\langle E_{\lambda}B_{T},E_{\lambda}B_{T}\rangle}$ (5)
$\leq\sqrt{\sum_{\lambda}\langle E_{\lambda}B_{S},E_{\lambda}B_{S}\rangle}\sqrt{\sum_{\lambda}\langle E_{\lambda}B_{T},E_{\lambda}B_{T}\rangle}$ (6)
$=\sqrt{\sum_{\lambda}\langle B_{S},B_{S}\rangle}\sqrt{\sum_{\lambda}\langle B_{T},B_{T}\rangle}$
$=\sqrt{\left\lvert S\right\rvert}\sqrt{\left\lvert T\right\rvert}$
$=\langle B_{T},B_{T}\rangle.$
Note that Inequaltiy (3) is tight if and only if all the nonzero products $f_{t}(\lambda)\langle E_{\lambda}B_{S},B_{T}\rangle$ have the same complex argument. Inequality (4) is tight if and only if $f_{t}(\lambda)=\pm 1$. Inequality (5) is tight if and only if $E_{\lambda}B_{S}$ is a scalar multiple of $E_{\lambda}B_{T}$. Inequality (6) is tight if and only if $\langle E_{\lambda}B_{S},B_{S}\rangle=\langle E_{\lambda}B_{T},B_{T}\rangle$. These motivate the following definitions.
@@
$\Lambda_{S}=\{\lambda:E_{\lambda}B_{S}\neq 0\}.$
We say two subsets $S$ and $T$ of vertices in $Y$ are strongly-cospectral if for each eigenvalue $\lambda$ of $H$,
$E_{\lambda}B_{S}=\pm E_{\lambda}B_{T}.$
If $S$ and $T$ are strongly-cospectral, then $\Lambda_{S}=\Lambda_{T}$, which can be partitioned into two disjoint sets
$\Lambda_{S,T}^{+}=\{\lambda:E_{\lambda}B_{S}=E_{\lambda}B_{T}\neq 0\}$
@@
$\Lambda_{S,T}^{-}=\{\lambda:E_{\lambda}B_{S}=-E_{\lambda}B_{T}\neq 0\}.$
@@
Let $W\subseteq\operatorname{col}(C_{a}+I)$ and $V\subseteq\operatorname{col}(C_{b}+I)$ be two subspaces with equal dimension. There is perfect $(W,V)$-transfer from $a$ to $b$ at time $t$ if and only if relative to some orthonormal coin basis induced by $(W,V)$, the Hermitian adjacency matrix $H$ satisfies the following.
@@
    $t\arccos\lambda$ is an even multiple of $\pi$ for each eigenvalue $\lambda\in\Lambda_{S,T}^{+}$.
@@
    $t\arccos\lambda$ is an odd multiple of $\pi$ for each eigenvalue $\lambda\in\Lambda_{S,T}^{-}$.
Proof. Suppose first that $f_{t}(H)B_{S}=B_{T}$. Then $\langle f_{t}(H)B_{S},B_{T}\rangle=\langle B_{T},B_{T}\rangle$. Thus $\langle f_{t}(H)B_{S},B_{T}\rangle>0$, and all inequalities in (3), (4), (5), (6) are tight. In particular, tightness of (3) now implies
$f_{t}(\lambda)\langle E_{\lambda}B_{S},B_{T}\rangle\geq 0,\quad\forall\lambda.$ (7)
Let $\lambda\in\Lambda_{S}$. Since equality holds in (5), there is a scalar $\alpha_{\lambda}$ such that
$E_{\lambda}B_{S}=\alpha_{\lambda}E_{\lambda}B_{T}.$
Plugging this into (7) shows
$\alpha_{\lambda}f_{t}(\lambda)>0.$
Since (6) is tight, we must have $\alpha_{\lambda}=\pm 1$. Finally, by tightness of (4), if $\alpha_{\lambda}=1$, then $f_{t}(\lambda)=1$, in which case $t\arccos\lambda$ is an even multiple of $\pi$, and if $\alpha_{\lambda}=-1$, then $f_{t}(\lambda)=-1$, in which case $t\arccos\lambda$ is an odd multiple $\pi$. This proves conditions (i) – (iii).
@@
$f_{t}(H)B_{S}$ $=\sum_{\lambda}f_{t}(\lambda)E_{\lambda}B_{S}$
$=\sum_{\lambda\in\Lambda_{S,T}^{+}}E_{\lambda}B_{S}-\sum_{\lambda\in\Lambda_{S,T}^{-}}E_{\lambda}B_{S}$
$=\sum_{\lambda\in\Lambda_{S,T}^{+}}E_{\lambda}B_{T}+\sum_{\lambda\in\Lambda_{S,T}^{-}}E_{\lambda}B_{T}$
$=\sum_{\lambda}E_{\lambda}B_{T}$
$=B_{T}.$
@@
We say $U$ is subspace periodic at $a$ if there is a time $t\in{\mathbb{R}}$ and a subspace $W$ of $\operatorname{col}(C_{a}+I)$ such that $x_{a}(W)$ is $U^{t}$-invariant; when $W$ is known, $U$ is said to be $W$-periodic at $a$. If, in addition,
$U^{t}x_{a}(w)=x_{a}(w)$
@@
Let $W$ be a subspace of $\operatorname{col}(C_{a}+I)$ with an orthonormal basis $\{w_{1},w_{2},\ldots,w_{d}\}$. Let $H$ be the Hermitian adjacency matrix relative to any orthonormal coin basis containing
$\{x_{a}(w_{1}),x_{a}(w_{2}),\cdots,x_{a}(w_{d})\}.$
Then $U$ is pointwise $W$-periodic at $a$ with period $t$ if and only if $t\arccos\lambda$ is an even multiple of $\pi$ for each $\lambda\in\Lambda_{S}$.
@@
[23, Ch 3] Given $\alpha_{1},\alpha_{2},\ldots,\alpha_{n}\in{\mathbb{R}}$ and $\beta_{1},\beta_{2},\ldots,\beta_{n}\in{\mathbb{R}}$, the following are equivalent.
@@
$\{\langle f_{t}(H)B_{S},B_{T}\rangle:t\in{\mathbb{R}}\}$
has $d$ as a limit point. It follows from the inequalities in (5) and (6) that $S$ and $T$ need to be strong-cospectrality to allow pretty good $(W,V)$-transfer. Given (i), set $\sigma_{\lambda}=0$ if $\lambda\in\Lambda_{S,T}^{+}$, and $\sigma_{\lambda}=1$ if $\lambda\in\Lambda_{S,T}^{-}$. Pretty good $(W,V)$-transfer occurs from $a$ to $b$ if and only if for any $\epsilon>0$, there is a $t\in{\mathbb{R}}$ such that
@@
5 Pointwise subspace state transfer
Sometimes we may want to preserve the amplitude on each arc during state transfer. We say there is pointwise perfect subspace state transfer from $a$ to $b$ if there is some time $t\in{\mathbb{R}}$ and some subspace $W\subseteq\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$ such that for any vector $w\in W$, there is a unimodular $\gamma$ such that
$U^{t}x_{a}(w)=\gamma x_{b}(w).$
Most studies on perfect and pretty good state transfer via weighted Grover coins fall into this category; the following is an example with $W=\mathrm{span}\{{\bf 1}\}$ and $\gamma=1$. When $W$ is known, this is called pointwise perfect $W$-transfer from $a$ to $b$. It follows from linearity of $U$ that if this happens, then the phase $\gamma$ does not depend on the choice of $w$.
$1111$
Figure 6: At time 0
$1111$
Figure 7: At time 66
As a strengthening of subspace state transfer, pointwise $W$-transfer from $a$ to $b$ imposes conditions on a special type of orthonormal coin bases, each of which contains
$\{x_{a}(w_{1}),x_{a}(w_{2}),\cdots,x_{a}(w_{d}),x_{b}(w_{1}),x_{b}(w_{2}),\cdots,x_{b}(w_{d})\},$
for some orthonormal basis $\{w_{1},w_{2},\ldots,w_{d}\}$ of $W$. We will call such a basis an orthonormal coin basis induced by $W$.
5.1 Lemma.
Let $W$ be a subspace of $\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$. Pointwise perfect $W$-transfer from $a$ to $b$ occurs at time $t$ if and only if there is a unimodular $\gamma\in{\mathbb{C}}$ such that
$f_{t}(H)B_{S}=\gamma B_{T}$
relative to some (or any) orthonormal coin basis induced by $W$.
Thus if $S$ and $T$ are the $W$-clones of $a$ and the $W$-clones of $b$, respectively, then a necessary condition for pointwsie $W$-transfer is that, for some unimodular $\gamma$,
$E_{\lambda}B_{S}=\pm\gamma E_{\lambda}B_{T}$
for each eigenprojection $E_{\lambda}$ of the special Hermitian weighted digraph. Following Section 4, we say $S$ is $\gamma$-strongly-cospectral to $T$ if the above holds, and define
$\Lambda_{S,T}^{\gamma}=\{\lambda:E_{\lambda}B_{S}=\gamma E_{\lambda}B_{T}\neq 0\}$
and
$\Lambda_{S,T}^{-\gamma}=\{\lambda:E_{\lambda}B_{S}=-\gamma E_{\lambda}B_{T}\neq 0\}.$
By relaxing the notion of strong-cospectrality, we obtain a characterization of pointwise $W$-transfer relative to an arbitrary orthonormal coin basis induced by $W$.
5.2 Theorem.
Let $W$ be a subspace of $\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$. Pointwise perfect $W$-transfer from $a$ to $b$ occurs at time $t$ if and only if there is a unimodular $\gamma\in{\mathbb{C}}$ such that relative to some (or any) orthonormal coin basis induced by $W$, the following hold.
  1. (i)
    $S$ is $\gamma$-strongly-cospectral to $T$.
  2. (ii)
    $t\arccos\lambda$ is an even multiple of $\pi$ for each eigenvalue $\lambda\in\Lambda_{S,T}^{\gamma}$.
  3. (iii)
    $t\arccos\lambda$ is an odd multiple of $\pi$ for each eigenvalue $\lambda\in\Lambda_{S,T}^{-\gamma}$.
Similarly, we say there is pointwise pretty good subspace state transfer from $a$ to $b$ if there is a subspace $W$ of $\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$ such that for any $\epsilon>0$ and any vector $w\in W$, there exist a time $t$ and a unimodular $\gamma$ such that
$\left\lVert U^{t}x_{a}(w)-\gamma x_{b}(w)\right\rVert<\epsilon.$
5.3 Theorem.
Let $W$ be a subspace of $\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$. Pointwise pretty good $W$-transfer occurs from $a$ to $b$ if and only if there is a unimodular $\gamma\in{\mathbb{C}}$ such that relative to some (or any) orthonormal coin basis induced by $W$, the following hold.
  1. (i)
    $S$ is $\gamma$-strongly-cospectral to $T$.
  2. (ii)
    $\sum_{\lambda\in\Lambda_{S}}\ell_{\lambda}\arccos\lambda=0,$ (9)
    we have
    $\sum_{\lambda\in\Lambda^{-\gamma}_{S,T}}\ell_{\lambda}\equiv 0\pmod{2}$.
Moreover, the target states can be approximated in integer steps if and only if (i) and (ii) hold with Equation (9) replaced by